Grade 12thPhysical Chemistry

How many electrons in an atom have the following quantam numbers? (a) n=4 ms=-1/2(b) n=3 l=0

Atharva koli
8 Years agoGrade 12th

1 Answer

Arun
8 Years ago
Dear Atharva
 
(a) Total number of electrons in an atom for a value of n = 2n∴ For n = 4, 
Total number of electrons = 2 (4) = 32 
The given element has a fully filled orbital as 1s2 2s2 2p6 3s2 3p6 4s2 3d10
Hence, all the electrons are paired. 
∴ Number of electrons (having n = 4 and m = - 1/2 ) = 16 
(b) n = 3, l = 0 indicates that the electrons are present in the 3s orbital. Therefore, the number of electrons having n = 3 and l = 0 is 2.
 
Regards
Arun (askIITians forum expert)